OpenStudy (anonymous):

What is the relative maximum and minimum of the function? f(x) = 2x^2 + 28x - 8 A. Minimum Value: -106 Range y > -7 B. Minimum Value: -106 Range y > -106 C. Minimum Value: 7 Range y > 7 D. Minimum Value: -7 Range y > -7

OpenStudy (campbell_st):

well since its a concave up parabola its going to have a minimum. do you know calculus...?

OpenStudy (anonymous):

Nope lol

OpenStudy (anonymous):

It only has minimum.

OpenStudy (campbell_st):

ok... well use the line of symmetry... to help find the vertex... this is where the minimum will occur the general form for the line of symmetry is \[x = \frac{-b}{2 \times a}\] in your question you have b = 28 and a = 2 can you calculate the value for the line of symmetry?

OpenStudy (anonymous):

x= -28/ 2*2? i can put that into a calculator and what i get will be my answer right?

OpenStudy (campbell_st):

thats right...but make sure the denominator is (2 x 2) in brackets

OpenStudy (anonymous):

okay :) Let me plug it in

OpenStudy (anonymous):

I got -7 ...but idk what the other number is ..

OpenStudy (anonymous):

Its either A or D

OpenStudy (campbell_st):

great... so plug x = -7 into your equation and that will give the minimum value of the curve.

OpenStudy (anonymous):

Oh okay lol

OpenStudy (campbell_st):

you need to take care with the substitution as 2*(-7)^2 is a positive number...

OpenStudy (anonymous):

98?

OpenStudy (campbell_st):

thats great so its then \[98 + 28 \times (-7) - 8\] to get the minimum value

OpenStudy (anonymous):

Its says math processing error ..

OpenStudy (anonymous):

Im thinking its A though :) So thanks for taking me step by step :)

OpenStudy (anonymous):

It's D @SkiTTleoooo47